OpenStudy (anonymous):

Find the side length with the given area 121in. PLEASE HELP!!!!

OpenStudy (anonymous):

side length of what?

OpenStudy (anonymous):

square

OpenStudy (anonymous):

sqrt(121) = 11.

OpenStudy (anonymous):

a square's sides are all congruent. to find the area of something, you multiply the length of the base by the length of the height. so let's say x is the length of the base of a square. since base = height: x*x = area x^2 = area we know the area x^2 = 121 solve for x: sqrt(x^2) = sqrt(121) |x| = 11

OpenStudy (anonymous):

Find the side length of a square with the given area 100 m2

OpenStudy (anonymous):

11. I'm pretty sure. @IAmSinged is correct.

OpenStudy (anonymous):

10m

OpenStudy (anonymous):

10. I would use the sqrt button on your calculator for these problems.

OpenStudy (anonymous):

evaluate the square root 256

OpenStudy (anonymous):

It's 16.

OpenStudy (anonymous):